the area of a square, expressed in square inches, is 165 in inches greater than its perimeter, expressed in inches. what is the length of a side of the square?
Answer by ankor@dixie-net.com(22740)

the area of a square, expressed in square inches, is 165 in inches greater than its perimeter, expressed in inches.
what is the length of a side of the square?
:
Area = x^2
Perimeter = 4x
:
x^2 - 4x = 165
:
x^2 - 4x - 165 = 0; our old friend, the quadratic equation!
Factor
(x - 15)(x + 11) = 0
Positive solution is what we want here
x = 15 is the length of the square
:
:
Is this true?
15^2 - 4(15) =
225 - 60 = 165